OpenStudy (anonymous):

Find the slope and the y-intercept: x+ y = 9

OpenStudy (anonymous):

\[\large y=\color{red}mx+\color{blue}b\]m is the slope b is the y-intercept Get the equation is this form

OpenStudy (anonymous):

So, we can easily answer this question by changing its form into y = mx + b. In this form, m is the slope and b is the y-intercept. x + y = 9 y = -x + 9 So, the slope is -1, and the y-intercept is 9

OpenStudy (phi):

you should have tried this one on your own, because all you have to do is add -x to both sides to change the equation into slope intercept form

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

so it would be: